OpenStudy (frame):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (anonymous):

it has to be the first equation because y-6 = -3(x+1) is the unsimplified form of the same equation that substitutes the coordinates into y - y1 = m(x - x1)